# On the PM2.5 – Mortality Relationship: A Bayesian Dynamic Model for Spatio-Temporal Confounding

### v1

@@ 6 Application to the Italian Mortality Data
We now apply the proposed DGLM to the Italian mortality dataset (see Section 3). The overall trend, $\widehat{X_{t}}$, is obtained using a natural cubic spline of time with $4$ degrees of freedom per year. For each HD and time instant, ${}^{m}{\bf Z}_{it}$ contains bases from a natural cubic spline of daily temperature and daily relative humidity (with 7 and 4 degrees of freedom, respectively), indicator variables for weekdays, for holidays, and for the lockdown periods.
@@
Figure 6 depicts the estimated local exposure effect, expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red solid line represents its posterior mean, while the shaded area the corresponding 95% credible interval (CI). The exposure effect is not constant in time, as no horizontal straight line can be drawn within the CI. This effect, instead, exhibits varying peaks across different years, and three phases are clearly identifiable: before (2018-19), during (2020-21) and after (2022) the pandemic outbreak.
During the first two years, the local exposure effect exhibits a seasonal pattern, with the peaks occurring during the summer months. The estimated association is bounded away from zero from July to October 2018 from April to September 2019, with the highest change in mortality risk of $7.47\%$ (95% CI: $[3.88\%,11.30\%]$) occurring on June 29, 2019.
Figure 6: Time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations estimated by the proposed DGLM. The red solid line represents its posterior mean, while the shaded area the corresponding 95% CI.
@@
These seasonal changes of the risk might be due to potential interaction effects of PM2.5 concentrations with air temperature. As noted in Section 1, increased ventilation and physiological susceptibility to air pollution may explain the higher percent change of risk during summer. Of course, interaction with other variables, such as other pollutants, are also possible (see, e.g., Liu et al., 2023, Ji et al., 2022, Siddika et al., 2019).
@@
During the pandemic phase, the CI always contains the zero except for few days in October 2020. The literature reveals a lack of consensus regarding the association between air pollution and adverse COVID-19 health outcomes in the general population (see, for example, Brunekreef et al., 2021a, Walton et al., 2021). More recently, the review by Bhaskar et al. (2023) suggests a positive association in the majority of previous studies. Our findings, instead, indicate that the local exposure effect in 2020 and 2021 is generally not bounded away from zero. To justify this, consider first of all that our model uses all-cause mortality in the elderly population as the health outcome, so it is not directly comparable to the COVID-19 studies mentioned earlier. Secondly, the issuance of the lockdown measures by the Italian government, as well as the mandatory use of facial masks, not only effectively curbed the transmission of the SARS-CoV-2 virus (and of other viruses), but also potentially reduced exposure to air pollution. During the lockdown periods, people were required to stay at home, resulting in reduced exposure to outdoor air pollutants. The ventilation was limited because the lockdowns occurred during non-summer months. Moreover, the use of facial masks may have reduced the inhalation of airborne hazardous particles (Kodros et al., 2021, Zhou et al., 2018).
@@
In the third phase, people started to return to their normal activities, resulting in increased exposure to air pollution. In fact, the local exposure effect has a shape similar to that in 2018, but is bounded away from zero only in July 2022 (probably because of mortality displacement), with the highest change in mortality risk of $4.82\%$ (95% CI: $[0.65\%,9.46\%]$) occurring on July 26.
@@
In Section G of the Supplementary Material we want to understand the consequences of removing the measured confounders (i.e., temperature, relative humidity, and indicators for weekdays, national holidays, and lockdown periods) from the model specification. Therefore, we impose the restriction that the elements of $\bm{\alpha}_{m}$ in Equation (6) are zero. The results are insightful, as the estimated exposure effect greatly overlaps the estimate obtained when the measured confounders are included, with differences occurring only in summer. This suggests that the proposed DGLM can provide accurate estimates even in complex scenarios, and that the confounding effect may be time-varying and may be due to heat waves.
@@ 7 Discussion
Motivated by a real case-study, in this paper we have proposed a DGLM as a solution for two types of bias that may affect the estimation and interpretation of the exposure–outcome relationship: misspecification and confounding bias. Misspecification bias arises from assuming that the exposure effect is constant in time when it is not. When analyzing the health effects of air pollution, Chiogna and Gaetan (2002) and Peng et al. (2005), among others, argue that there is uncertainty about the “true” functional form of the exposure–outcome association. Therefore, assuming a constant effect may be too restrictive and result in biased estimates. We acknowledge this fact and assume that there exists a smooth (but unknown) function of the exposure, $f(\cdot;\cdot)$, describing the association of interest. Confounding bias is due to the presence of an unmeasured confounder. To control for it, we decompose the exposure effect into large- and small-scale components following Janes et al. (2007), but, differently from them, we assume that the local effect is time-varying.
@@
We have shown that the proposed approach enjoys some advantages. On one hand, it is able to capture any time-varying exposure effect, thus overcoming the misspecification bias. On the other hand, thanks to the first-order Taylor’s expansion, the DGLM can model the local exposure effect within a flexible framework, thus mitigating the confounding bias. In this sense, the proposed approach constitutes an extension of the model suggested by Janes et al. (2007).
@@
The effectiveness of the proposed approach has been demonstrated through artificial data analyses. We have shown that it performs better than the Null and the GLMadj models. In particular, the proposed DGLM not only can reduce the bias, but also can capture complex temporal fluctuations of the exposure effect and almost recover its true functional form. The bias is less alleviated when the unmeasured variable is less smooth than the exposure in time, but it is worth noticing that this situation is rather unrealistic in practice. We have seen that the GLMadj model could achieve the lowest average mean squared error in the case of a constant exposure effect (see Table E.4), however, the shape of the exposure–outcome association is typically unknown. Unless there is strong prior information about this shape, we thus suggest using the proposed DGLM. Furthermore, as a proof of concept, in Section F of the Supplementary Material, we set up a scenario where a complete absence of confounders is assumed. In this case, the exposure effect results fully recovered.
@@
Many previous empirical works assume an exposure–outcome association constant in time. However, the evidence from the analysis of the Italian mortality data have indicated that this assumption is conceivably incorrect. In fact, the estimated local exposure effect exhibits a seasonal pattern, with a higher percent change of risk observed during the summer months, when the overall PM2.5 level is lower. Similar conclusions were reached in other studies (Peng et al., 2005, 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018). Although this fact is apparently counterintuitive, it has been observed that it might be due to possible interaction effects of PM2.5 concentrations with meteorological variables, such as temperature, and other air pollutants. We have also noted that our result may be related to possible effect modification due to changes in the particulate matter mixture (Franklin et al., 2008, Hao et al., 2023, Masselot et al., 2022). The relative proportions of the PM2.5 components have not been analyzed in this study, so they could be the subject of future research. However, it is important to note that Hao et al. (2023) found a non-linear long-term association between some PM2.5 constituents and mortality.
@@
An exception to the seasonal behavior are the years 2020 and 2021, when the exposure effect is not bounded away from zero. This can be justified by the fact that the lockdown measures, such as stay-at-home orders and the use of masks, have likely reduced the exposure to air pollution. It is worth recalling that none of the competing approaches in Section 2 were able to capture this behavior. Additionally, our findings may contradict the results of the majority of studies reviewed in Bhaskar et al. (2023). However, there are differences in the definition of outcomes between our analysis and theirs, as we focused on all-cause mortality in the elderly population. Moreover, not all studies related to COVID-19 properly account for potential confounders, so their findings should be interpreted with caution (Bhaskar et al., 2023, Brunekreef et al., 2021b, Walton et al., 2021). In contrast, the proposed DGLM is able to almost completely capture the exposure effect even in the absence of measured confounders (see Figure G.5 in the Supplementary Material).
To account for the residual spatial autocorrelation in the outcome, possibly due to unmeasured confounding, the proposed model can potentially provide for the adjacency of areal units, a feature often ignored in multi-site time-series studies (see, e.g., Dai et al., 2014, Janes et al., 2007, Peng et al., 2005, Qiu et al., 2022, Zhou et al., 2021). In fact, this could be achieved by including into the linear predictor a spatially structured random effect with a conditional autoregressive specification. On the other hand, our model cannot consider spatial variations of the local exposure effect, so it would be interesting to specify a spatially-varying coefficients model, which would allow discovering differences due to urbanicity and morphology of the study area, for example.
@@
It is arguable that, even if the proposed DGLM could recover the true exposure effect, $f(\cdot;\cdot)$, its functional form still remains unknown. In other words, the complete list of interaction terms cannot be determined. However, we believe that this is not a drawback only of the proposed model. In fact, the list of interaction terms is unknown when fitting any regression model, and it is up to the researcher to choose the most appropriate form for $f(\cdot;\cdot)$. This decision may involve some degree of uncertainty. The proposed approach, instead, allows for more flexibility and should be able to capture the effect of lockdowns as well as interaction effects between the exposure and other variables, such as other pollutants, so that inferential problems due to misspecification are overcome.
In conclusion, the proposed DGLM is a flexible and effective approach to estimate the exposure–outcome association in the presence of time-varying effects and unmeasured spatio-temporal confounding. The model is able to capture the complex dynamics of the exposure effect, and to control for confounding. Also, in the scenario simulated without confounding, the true effect is correctly recovered. The proposed approach is a valuable tool for researchers interested in the health effects of air pollution, and it can be extended to other fields where the exposure–outcome association is of interest.

### v2

@@ 7 Application to the Italian Mortality Data
We apply SDGLMC to the Italian mortality dataset (see Section 2), and compare its results to those from the competing approaches. The large-scale trend, $\widehat{X}_{it}$, is obtained using a principal spline with $5$ spatial bases and $20$ temporal bases. For each HD and time instant, ${\bf M}_{it}$ contains bases from a natural cubic spline of daily temperature and daily relative humidity (with 7 and 4 degrees of freedom, respectively), and indicator variables for weekdays, holidays and lockdown periods (hence ${p_{m}}=19$; see also Section A of the Supplementary Material).
Types of Space-Time Interaction Component, $\delta_{0it}^{*}$
We can envision several distributional specifications for the space-time interaction component of the intercept, $\delta_{0it}^{*}$, depending on the correlation structure we want to impose. Knorr-Held (2000) proposes four types of models, giving rise to four types of space-time interaction terms, characterized by different prior assumptions. These are reviewed by Franco-Villoria et al. (2022) and Orozco-Acosta et al. (2023), and are summarized in Table 2. Type T1 assumes unstructured space-time variability, and is obtained by setting $\delta_{0it}^{*}=0$ for all $i$ and $t$, since an unstructured error is already present in Equation (7). Type T2 assumes independent temporal trends for each HD, with a common variance for the innovations. Type T3 assumes independent spatial trends for each time instant, with a common conditional variance. Type T4 assumes complete dependence, i.e. the temporal trends are different from HD to HD, but are more likely to be similar if the HDs are neighbors.
The prior assumptions discussed in Knorr-Held (2000) can be further relaxed. In fact, during the pandemic, some HDs in Lombardy – more specifically in the province of Bergamo – experienced a much greater increase in deaths than any other HD. Therefore, a fifth interaction type, T5, is also considered, assuming independent space-time interactions across the HDs that follow random walk processes with area-specific evolution variances. This captures temporal trends that differ considerably among HDs but do not have any spatial structure.
Table 2: Space-time interaction types considered for $\delta_{0it}^{*}$. DIC values for the corresponding models are reported, with the lowest value highlighted in italics.
Type Space-Time structure Distributional Assumption DIC
T1 Absent $\delta_{0it}^{*}=0$ 814038
T2 Independent RW, same variance $\delta_{0it}^{*}\sim\mathcal{N}(\delta_{0i,t-1}^{*},\sigma_{w0}^{*2})$ 807711
T3 Independent ICAR $\bm{\delta}_{0t}^{*}\sim ICAR(\sigma_{w0}^{*2})$ 811293
T4 Complete dependence $\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*}\sim ICAR(\sigma_{w0}^{*2})$ 811923
T5 Independent RW, different variances $\delta_{0it}^{*}\sim\mathcal{N}(\delta_{0i,t-1}^{*},\sigma_{w0i}^{*2})$ 803861
Model Fitting and Presentation of Results
@@
After fitting a model for each of the five types of space-time interaction components, the one with the lowest deviance information criterion (DIC; Spiegelhalter et al., 2002) is selected. Based on the last column of Table 2, the model with interaction type T5 has the smallest DIC value, hence its results will be discussed below. A reason for this may be that, in comparison to the models proposed by Knorr-Held (2000), the model with T5 seems more able to capture very specific spatio-temporal trends, such as those observed in the Bergamo area due to the pandemic.
@@
Figure 7 depicts the estimated space-averaged local exposure effect (i.e., $\overline{\delta}_{1}+\delta_{1t}$), expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red solid line represents its posterior mean, while the shaded area the corresponding 95% credible interval (CI). The exposure effect is not constant in time, as no horizontal straight line can be drawn within the CI. This effect, instead, exhibits varying peaks across different years, and three phases are clearly identifiable: before (2018-19), during (2020-21) and after (2022) the pandemic outbreak.
During the first two years, the local exposure effect exhibits a seasonal pattern, with the peaks occurring during the summer months. The estimated association is bounded away from zero from August to September 2018 from April to October 2019, with the highest change in mortality risk of $5.50\%$ (95% CI: $[3.03\%,7.98\%]$) occurring on June 29, 2019.
Figure 7: Time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations estimated by SDGLMC. The red solid line represents its posterior mean, while the shaded area the corresponding 95% CI.
@@
These seasonal changes of the risk suggest that the unmeasured confounders could have not only a marginal effect on mortality (hence influencing $\beta_{0it}$), but also an interaction effect with the exposure (hence affecting $\beta_{1it}$). In fact, the seasonal pattern might be due to potential interactions between PM2.5 concentrations and air temperature. As noted in Section 1, increased ventilation and physiological susceptibility to air pollution may explain the higher percent change of risk during summer. Besides, interaction with other variables, such as other pollutants, are also possible (see, e.g., Liu et al., 2023, Ji et al., 2022, Siddika et al., 2019).
@@
During the pandemic phase, the CI always contains the zero except for few days during December 2021. The literature reveals a lack of consensus regarding the association between air pollution and adverse COVID-19 health outcomes in the general population (see, for example, Brunekreef et al., 2021a, Walton et al., 2021). More recently, the review by Bhaskar et al. (2023) suggests a positive association in the majority of previous studies. Our findings, instead, indicate that the local exposure effect in 2020 and 2021 is generally not bounded away from zero (except for few days). To justify this, consider first of all that our model uses all-cause mortality in the elderly population as the health outcome, so it is not directly comparable to the COVID-19 studies mentioned earlier. Secondly, the issuance of the lockdown measures by the Italian government, as well as the mandatory use of facial masks, not only effectively curbed the transmission of the SARS-CoV-2 virus (and of other viruses), but also potentially reduced exposure to air pollution. During the lockdown periods, people were required to stay at home, resulting in reduced exposure to outdoor air pollutants. The ventilation was limited because the lockdowns occurred mainly during winter months. Moreover, the use of facial masks may have reduced the inhalation of airborne hazardous particles (Kodros et al., 2021, Zhou et al., 2018).
@@
During the third phase, people started to return to their normal activities, resulting in increased exposure to air pollution. In fact, the local exposure effect has a shape similar to that in the first phase, and is bounded away from zero from July to September 2022, with the highest change in mortality risk of $3.53\%$ (95% CI: $[0.61\%,6.44\%]$) occurring on July 26.
Comparison with Competing Approaches
For each approach introduced in Section 5, Figure 8 shows the pooled estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 levels. The posterior mean from SDGLMC is also shown (in red) for comparison. The Null model yields the highest estimate on average ($3.92\%$), whereas the GLMadj model estimates the lowest percent change on average ($0.21\%$). The models (i)–(iii) assume that the association of interest is constant over time, however we have argued that this assumption is unlikely.
Figure 8: Estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations: comparison of SDGLMC with the approaches described in Section 5.
The only two approaches able to capture time-varying strengths of the association are the Dummy and the Periodic models. They seem to agree with SDGLMC on the temporal evolution of the effect, showing peaks in summer and troughs in winter. The Dummy and the Periodic models, however, have some limitations, since both assume that the effect changes within the year, but is the same across years. This is not consistent with the evidence from the data, as the pandemic outbreak may have altered the strength of the association.
In terms of spatial variation, we analyze the differences between area-specific and pooled estimates of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. Figure 9 shows the posterior means of those differences under each competing approach. It also shows the posterior mean of the spatial component of the local exposure effect, i.e. $\delta_{1i}$, expressed in terms of percent change of risk. The spatial variability of the exposure effect approaches zero in all cases except for the Null and GLMadj models. However, when considering the associated uncertainty, there is no evidence of spatial variation in the exposure effect for any of the models. This may be related to the relatively small size of the study region and the similarity between Piemonte and Lombardia (e.g., in terms of socio-economic indicators).
Figure 9: Spatial component of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations under each approach.
@@
In Section G of the Supplementary Material we want to understand the consequences of removing the measured confounders (i.e., temperature, relative humidity, and indicators for weekdays, national holidays, and lockdown periods) from the model specification. Therefore, we impose the restriction that the elements of $\bm{\alpha}$ in Equation (6) are zero. The results are insightful, as the estimated exposure effect greatly overlaps the estimate obtained when the measured confounders are included, with differences occurring only in summer. This suggests that SDGLMC can provide accurate estimates even in complex scenarios, and that the confounding effect may be time-varying and may be due to heat waves.
@@ 8 Discussion
Motivated by a real case-study, in this paper we have proposed the spatial dynamic generalized linear model for confounding (SDGLMC) as a solution for two types of bias that may affect the estimation and interpretation of the exposure–outcome relationship: misspecification and confounding bias. Misspecification bias arises from assuming that the exposure effect is constant in time when it is not. When analyzing the health effects of air pollution, Chiogna and Gaetan (2002) and Peng et al. (2005), among others, argue that there is uncertainty about the “true” functional form of the exposure–outcome association. Therefore, assuming a constant effect may be too restrictive and result in biased estimates. We acknowledge this fact and assume that there exists a smooth (but unknown) function of the exposure, $f(\cdot;\cdot)$, describing the association of interest. Confounding bias is due to concurvity. To control for it, we decompose the exposure effect into large- and small-scale components, then we want to recover the local exposure effect. The same objective is pursued by Janes et al. (2007) and Dupont et al. (2022), but, differently from these authors, we assume that the local effect is spacetime-varying. Hence, SDGLMC constitutes an extension of their models. Importantly, another connection exists between SDGLMC and the work of Dupont et al. (2022), as the smooth function of spatial coordinates in their model and the spatio-temporal intercepts in SDGLMC play a similar role.
@@
We have shown that SDGLMC enjoys some advantages. On one hand, it is able to capture any spacetime-varying exposure effects, thus overcoming the misspecification bias. On the other hand, thanks to the first-order Taylor’s expansion, SDGLMC can model the local exposure effect within a flexible framework. To recover the local, short-term association, the outcome is regressed on a detrended exposure. This is useful because detrending improves estimation stability and helps reduce the confounding bias (Janes et al., 2007, Greven et al., 2011).
@@
The effectiveness of the proposed approach has been demonstrated through artificial data analyses. We have shown that it performs better than several competitors. In particular, SDGLMC not only can reduce the bias, but also can capture complex spatio-temporal dynamics of the exposure effect and almost recover its true functional form. The bias is less alleviated when the unmeasured variable is less smooth than the exposure in time, but it is worth noticing that this situation is rather unrealistic in practice. Note that the models assuming a constant association (e.g., GLMadj) could achieve better performances if the true effect is constant, however the shape of the exposure–outcome association is typically unknown. Unless there is strong prior information about this shape, we thus suggest using SDGLMC.
@@
Many previous empirical works assume a constant exposure–outcome association. However, the evidence from the analysis of the Italian mortality data have indicated that this assumption is conceivably incorrect. In fact, the estimated local exposure effect exhibits a seasonal pattern, with a higher percent change of risk observed during the summer months, when the overall PM2.5 level is lower. Similar conclusions were reached in other studies (Peng et al., 2005, 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018). Although this fact is apparently counterintuitive, it has been observed that it might be due to possible interaction effects of PM2.5 concentrations with meteorological variables, such as temperature, and other air pollutants. We have also noted that our result may be related to possible effect modification due to changes in the particulate matter mixture (Franklin et al., 2008, Hao et al., 2023, Masselot et al., 2022). The relative proportions of the PM2.5 components have not been analyzed in this study, so they could be the subject of future research. However, it is important to note that Hao et al. (2023) found a non-linear long-term association between some PM2.5 constituents and mortality.
A reviewer pointed out that if the effect modification is solely due to measured confounders, it can be studied more straightforwardly by including interaction terms between PM2.5 and the confounders in the linear predictor. We explore this possibility in Section G.2 of the Supplementary Material. However, since the estimated effect is significantly different from that in Figure 7 (especially during the summer months and the pandemic), we conclude that the effect modification is due to both measured and unmeasured confounders.
@@
An exception to the seasonal behavior are the years 2020 and 2021, when the exposure effect is not bounded away from zero. This can be justified by the fact that the lockdown measures, such as stay-at-home orders and the use of masks, have likely reduced the exposure to air pollution. A similar result is found in the absence of measured confounders (see Figure G.1 in the Supplementary Material), and is also suggested by recent studies (see, e.g., Hadei et al., 2021, López-Feldman et al., 2021). It is worth recalling that none of the competing approaches in Section 2 were able to capture this behavior. However, separate analyses per year (see Table 3) suggest patterns that are qualitatively similar to our results. We limited this analysis to the GLMadj and JDZ models due to the limited temporal extent. Nevertheless, both models indicate a non-significant exposure effect during the second phase.
Table 3: Posterior means obtained from fitting the GLMadj and JDZ approaches separately for each year. The values are expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The 95% CIs are reported within the square brackets.
2018 2019 2020 2021 2022
GLMadj 0.75 [0.06, 1.45] 0.81 [0.22, 1.39] 0.14 [-0.43, 0.69] -0.03 [-0.69, 0.68] 0.67 [0.01, 1.31]
JDZ 0.82 [0.13, 1.50] 0.85 [0.25, 1.43] 0.26 [-0.30, 0.81] -0.03 [-0.69, 0.68] 0.58 [-0.07, 1.22]
To account for the residual spatial autocorrelation in the outcome, possibly due to unmeasured confounding, SDGLMC accounts for the adjacency of areal units. That is, spatial and space-time components of the intercept are included into the linear predictor. Besides, SDGLMC allows to possibly capture the spatial variations of the local exposure effect, which could potentially discover differences due to urbanicity and morphology of the study area, for example. However, in the application to the Italian mortality data, the local exposure effect is homogeneous across space, probably because the study area is relatively small, and Piemonte and Lombardia have similar socio-economic and environmental characteristics.
@@
It is arguable that, even if SDGLMC could recover the true exposure effect, $f(\cdot;\cdot)$, its functional form still remains unknown. In other words, the complete list of interaction terms cannot be determined. However, we believe that this is not a drawback only of the proposed model. In fact, the list of interaction terms is unknown when fitting any regression model, and it is up to the researcher to choose the most appropriate form for $f(\cdot;\cdot)$. This decision may involve some degree of uncertainty. The proposed approach, instead, allows for more flexibility and should be able to capture the effect of lockdowns as well as interaction effects between the exposure and other variables, such as other pollutants, so that inferential problems due to misspecification are overcome.
Further research could focus on the causal interpretation of the estimated exposure effect, in which case other assumptions are needed (see, e.g., Hernán and Robins, 2020). However, note that the analytical approach pursued here should not be considered as less credible than causal inference methods (Dominici and Zigler, 2017).
In conclusion, SDGLMC is a flexible and effective approach to estimate the exposure–outcome association in the presence of spacetime-varying effects and unmeasured spatio-temporal confounding. The model is able to capture the complex dynamics of the exposure effect, and to control for confounding. The proposed approach is a valuable tool for researchers interested in the health effects of air pollution, and it can be extended to other fields where the exposure–outcome association is of interest.